# Machine Learning Model Integration with Open World Temporal Logic for Process Automation

### v1

@@ Abstract
Recent advancements in Machine Learning (ML) have yielded powerful models capable of extracting structured information from diverse and complex data sources. However, a significant challenge lies in translating these perceptual or extractive outputs into actionable, reasoned decisions within complex operational workflows. To address these challenges, this paper introduces a novel approach that integrates the outputs from various machine learning models directly with the PyReason framework, an open-world temporal logic programming reasoning engine. PyReason’s foundation in generalized annotated logic allows for the seamless incorporation of real-valued outputs (e.g., probabilities, confidence scores) from diverse ML models, treating them as truth intervals within its logical framework. Crucially, PyReason provides mechanisms, implemented in Python, to continuously poll ML model outputs, convert them into logical facts, and dynamically recompute the minimal model, ensuring real-tine adaptive decision-making. Furthermore, its native support for temporal reasoning, knowledge graph integration, and fully explainable interface traces enables sophisticated analysis over time-sensitive process data and existing organizational knowledge. By combining the strengths of perception and extraction from ML models with the logical deduction and transparency of PyReason, we aim to create a powerful system for automating complex processes. This integration finds utility across numerous domains, including manufacturing, healthcare, and business operations.
@@ 1 Introduction
Recent advancements in Machine Learning (ML) have yielded powerful models capable of extracting structured information from diverse and complex data sources. Systems ranging from object detectors and image classifiers in computer vision to large language models (LLMs) in natural language processing can now identify patterns, classify inputs, and extract entities with high efficacy. These models generate valuable outputs–such as object locations with confidence scores, image classifications with probabilities, or textual intents with associated sentiments–that hold immense potential for automating and optimizing real-world processes.
However, a significant challenge lies in translating these perceptual or extractive outputs into actionable, reasoned decisions within complex operational workflows. ML models, while proficient in their specific tasks, often lack the inherent mechanisms for deep logical reasoning, transparently managing uncertainty, integrating explicit domain knowledge, or handling the temporal dynamics crucial for many automation scenarios. Bridging this gap requires a framework that can synthesize these varied ML outputs and subject them to rigorous, explainable logical analysis.
@@
To address these challenges, this paper introduces a novel approach that integrates the outputs from various machine learning models directly with the PyReason framework, an open-world temporal logic reasoning engine. Figure 1 provides a high-level overview of this integration framework. PyReason’s foundation in generalized annotated logic [8] allows for the seamless incorporation of real-valued outputs (e.g., probabilities, confidence scores) from diverse ML models, treating them as truth intervals within its logical framework. We note that existing work focuses on integration between machine learning models and logic-based reasoning by either modifying the ML model (e.g., see [5, 11, 12, 6, 7, 4]) or by adding rules about the failure of the model to the logic program (e.g., see [9, 2, 10]). In this paper, we examine practical methods complementary to both approaches - focusing on software engineering questions concerning the integration between logic programs and ML models.
@@
A central pillar of our proposal is PyReason’s ability to seamlessly interconnect logical programs with Python code. This is achieved through mechanisms that enable Python code to continuously poll ML model results, converting these perceptions into logical facts that are directly injected into the logic program. This dynamic process not only feeds the reasoning engine but also triggers the exact re-computation of the minimal model, ensuring that logical decisions always reflect the latest observations from the ML model. Given the ICLP audience’s familiarity with both Python and logic programming, we have prioritized a design that explicitly intersperses both paradigms, facilitating a direct understanding of the interaction between ML perception and logical deduction. We have made significant design decisions to simplify this integration and enhance usability. For instance, the introduction of specific base classes such as LogicIntegrationBase and TemporalLogicIntegratedClassifier standardizes the inference flow and the conversion of predictions into logical facts. These decisions not only simplify the syntax and the management of optional arguments, but also structure the interaction, allowing for a robust and modular solution for complex process automation.
In addition to this integration, PyReason offers native support for temporal reasoning, knowledge graph integration, and fully explainable inference traces, enabling sophisticated analysis over time-sensitive process data and existing organizational knowledge. By combining the strengths of perception and extraction from ML models with the logical deduction and transparency of PyReason, our objective is to create a powerful system for automating complex processes. This integration finds utility in numerous domains, including manufacturing (leveraging object detection for quality control), healthcare (using image classification for diagnostic support), and business operations (employing LLMs for document processing and customer interaction)
@@
The remainder of this paper is structured as follows: Section 2 introduces the PyReason framework, its logical foundations, and reasoning capabilities. Section 3 details the integration with machine learning models via a standardized interface. Section 4 presents two case studies–robotic welding and a card game–to illustrate practical applications. Section 5 concludes with a discussion of the broader implications.
@@ 2 The PyReason Framework
PyReason is a software framework developed for open-world temporal logic, implemented in Python. It is founded on generalized annotated logic, a system which not only encapsulates the current range of differentiable (i.e., fuzzy) first-order logics but also integrates temporal extensions to support inference over finite durations. A primary design objective of PyReason is to offer direct support for reasoning over graphical structures, such as knowledge graphs, social networks, and biological networks. The framework is engineered to generate fully explainable traces of its inference processes, ensuring transparency. Additionally, it incorporates practical functionalities like type checking and a memory-efficient implementation to conduct exact yet scalable deductive inference. PyReason endeavors to supply a modular and extensible platform for deduction, addressing the constraints of bespoke software frequently developed for particular logics within neuro-symbolic frameworks. It captures a breadth of capabilities observed in diverse neuro-symbolic frameworks, encompassing fuzzy logic, open-world reasoning, temporal reasoning, and graph-based reasoning. The PyReason software is publicly accessible via GitHub (https://github.com/lab-v2/pyreason).
@@ 2.1 Core Logical Principles and Reasoning over Graphs
The logical underpinning of PyReason is generalized annotated logic, augmented with various extensions to support a wide array of capabilities. A defining characteristic is its application of real-valued interval annotations, allowing logical statements to be annotated with elements from a lower semilattice structure composed of intervals that are subsets of $[0,1]$. This design aligns with truth intervals in fuzzy operators and paradigms in neuro-symbolic reasoning. Classical logic is supported with annotations $[0,0]$ (false) and $[1,1]$ (true), while tri-valued logic can use $[0,1]$ for ”unknown”. Negation of an atom annotated with $[l,u]$ results in an annotation of $[1-u,1-l]$ for its strong negation.
PyReason integrates temporal logic over finite periods through ”interpretations,” which are an initial set of facts correct at $t=0$. Temporal annotations $I(A,\hat{t}):[L,U]$ represent discrete time-points, where $\hat{t}$ is a specific time $t$ or $s$ for static interpretations (unchanging over time). Non-static predicates can change annotations over time without a monotonicity requirement. The framework operates under an open-world assumption: interpretations not explicitly defined are initially ”unknown” (e.g., $[0,1]$).
The framework is explicitly designed for reasoning over graphical structures like knowledge graphs. It assumes a graph $G=(\mathcal{C},E)$, where nodes $\mathcal{C}$ are constants and edges $E$ specify potential relationships. Predicates are typically unary (node labels) or binary (edge labels). A reserved binary predicate ‘rel‘ exists, where $rel(a,b)$ is a tautology if $(a,b)\in E$ and uncertain otherwise. Graphical syntactic extensions allow for operators like requiring the existence of $k$ items.
@@ 2.2 Key Reasoning Capabilities and Features
Logical rules are the primary mechanism for modifying annotations of non-static predicates. PyReason supports ground rules, universally quantified non-ground rules (for reasoning within or across edges), and rules with rule-based quantifiers in the head. Rules include a time delay $\Delta t$, specifying when the head is affected after the body is satisfied. Rule heads often involve functions $f(x_{1},...,x_{n})$ (e.g., T-norms, T-conorms, algebraic functions) to combine annotations from body literals.
@@
PyReason detects and manages logical inconsistencies. Inconsistency can occur if a new interpretation $[L^{\prime},U^{\prime}]$ is not a subset of the current $[L,U]$, or between an atom and its negation, or complementary predicates. PyReason flags and can report these inconsistencies. The explainable fixpoint operator allows tracing the cause. An option exists to resolve inconsistency by resetting the annotation to $[0,1]$ (uncertainty) and setting the atom to static.
@@ 2.3 Implementation and Neuro-Symbolic Relevance
PyReason is actualized as a contemporary Python-based framework, meticulously designed for scalability and correctness. Graphical input is via GraphML or a Networkx graph, with initial conditions and rules in text format. The Numba JIT compiler optimizes key operations for performance, and CPU parallelism is supported. For memory efficiency, only current interpretations are stored; past states are reconstructible via rule traces, which also ensure explainability. This design supports canonical and non-canonical models with minimal computational overhead by recomputing interpretations at each step.
PyReason is strategically positioned as an instrument in neuro-symbolic AI. Generalized annotated logic, its foundation, captures a wide spectrum of real-valued, temporal, and fuzzy logics used in frameworks like Logical Tensor Networks (LTN) and Logical Neural Networks (LNN), both viewable as subsets of annotated logic. PyReason can conduct inference on logic programs learned or modified by such systems. It furnishes crucial capabilities like graph-based and temporal reasoning, potentially absent in other neuro-symbolic frameworks’ native logics. The explainability of its inference is paramount for understanding and debugging complex neuro-symbolic models. By supporting generalized annotated logic and its extensions, PyReason facilitates system design independent of the learning process, offering a robust deductive engine.
@@ 3 Integration with Machine Learning Models
PyReason allows users to seamlessly integrate machine learning (ML) models within the logical framework. ML models output tensors are converted into PyReason facts which can be used to perform inference. This section outlines this process and illustrates its simplicity within our framework.
@@ 3.1 Logic Integrated Base Class
Pyreason defines an abstract base class, LogicIntegrationBase, designed for subclassing to integrate specific machine learning models. This class declares three protected methods that are called sequentially and must be implemented by every subclass:
@@ 3.2 Logic Integrated Classifier
For brevity and simplicity we outline the process of converting probabilities to annotated bounds as a flowchart depicted in Figure 2. By following this logic, we’re able to convert model output tensors into annotated facts which can be added into PyReason for downstream inference.
@@ 3.3 Automated Temporal Logic Integrated Classifier
In many real‐world applications, data arrive continuously from sensors or cameras, and logical inferences must adapt in real time. To address this, PyReason provides the TemporalLogicIntegratedClassifier, a subclass of LogicIntegratedClassifier that ties PyReason’s internal temporal reasoning to wall‐clock time. This class adds two key parameters: poll_interval, which specifies the frequency (in seconds or PyReason time) to poll the model with new input data if the poll_condition is either satisfied or not defined; and poll_condition, a PyReason query that determines whether a new inference cycle should be triggered.
At each interval, PyReason evaluates the poll_condit_ion. If the condition is satisfied–or if no condition is specified–the system automatically:
@@
This end-to-end loop (Figure 3) requires no additional glue code: PyReason handles the timing, model polling, fact injection, and inference in a single, explainable pipeline.
Section 4 illustrate two concrete workflows built on this machinery: defect detection in robotic welding and decision-making in a simulated card game. In each case, the temporal classifier ensures that as soon as new data are available, the logic program reflects the latest observations and derives the appropriate conclusions.
@@ 4.1 Welding Classifier
In an industrial robotic welding cell, detecting weld defects in real time is crucial to maintaining product quality and avoiding costly rework. Here, we apply a pre-trained weld-defect detector to identify gaps in each weld image, feed those results into PyReason to attempt automated gap closure, and mark the part as defective if the gap cannot be closed.
@@ 4.1.1 Logical Encoding in PyReason
The gap(weld_object) atom is added to the program by PyReason’s LogicIntegratedClassifier. These two rules help determine if the weld was unsuccessful and if after a repair was attempted there is still a gap, then the part is marked as defective. A flowchart of the pipeline can be seen in Figure 4. The repair and conclusion that a part might be defective happens in a background routine (Thread 2) while the main thread (Thread 1) goes over each object to be welded and performs the welding task.
We use the TemporalLogicIntegratedClassifier defined in Section 3.3, with a poll_interval set to 0.5 seconds and the poll_condition set to gap(weld_object):[1,1]. This allows us to check whether there is a gap in the weld every 0.5 seconds, and attempt to repair the gap in the background.
@@ 4.1.2 Interpreting the Inference Trace
Table 1 showcases a PyReason explainable trace after welding 5 objects over 7 fixed-point operations (FPO). It can be seen how both thread 1 and thread 2 update the interpretation simultaneously. Thread 1 is always called first on every new object and thread 2 is invoked only when a repair is necessary. In FPOs 2-3 a repair is attempted due to a gap being detected in FPO 1 and the object is deemed defective after an unsuccessful repair. Similary, another repair is invoked in FPO 5, but the repair is successful and the part is not marked as defective.
@@ 4.2 Simple Card Game
This case study demonstrates how PyReason’s open-world temporal logic can drive decision-making in a simple ”42” card game by periodically integrating outputs from an image classifier and applying annotated logic to manage uncertainty over time. The objective of the game is to draw as many cards as possible without exceeding a point total of 42. Card values are defined as 3 points for an Ace, 6 for numeric cards, and 9 for face cards. At each turn, a decision is made to draw another card if there is a non-zero probability of remaining at or below the 42-point limit, based on the cards remaining in the deck. The game concludes when any subsequent draw would guarantee exceeding tthe limit.

### v2

@@ Abstract
Recent advances in Machine Learning (ML) have produced models that extract structured information from complex data. However, a significant challenge lies in translating these perceptual or extractive outputs into actionable and explainable decisions within complex operational workflows. To address these challenges, this paper introduces a novel approach that integrates the outputs of various machine learning models directly with the PyReason framework, an open-world temporal logic programming reasoning engine. PyReason’s foundation in generalized annotated logic allows for the incorporation of real-valued outputs (e.g., probabilities, confidence scores) from a diverse set of ML models, treating them as truth intervals within its logical framework. Crucially, PyReason provides mechanisms, implemented in Python, to continuously poll ML model outputs, convert them into logical facts, and dynamically recompute the minimal model to enable decision-making in real-time. Furthermore, its native support for temporal reasoning, knowledge graph integration, and fully explainable interface traces enables an analysis of time-sensitive process data and existing organizational knowledge. By combining the strengths of perception and extraction from ML models with the logical deduction and transparency of PyReason, we aim to create a powerful system for automating complex processes. This integration is well suited for use cases in numerous domains, including manufacturing, healthcare, and business operations.
@@ 1 Introduction
Recent advances in Machine Learning (ML) have produced models that extract structured information from complex data. Object detectors, image classifiers, and large language models can now recognize patterns, classify inputs, and extract entities. Their outputs, such as object locations, image labels, or text sentiments, can help automate and improve real-world processes.
However, a significant challenge lies in translating these perceptual or extractive outputs into actionable and explainable decisions within complex operational workflows. ML models, while proficient in their specific tasks, often lack the inherent mechanisms for explainable logical reasoning, transparently managing uncertainty, integrating explicit domain knowledge, or handling the temporal dynamics crucial for many automation scenarios. Bridging this gap requires a framework that can synthesize these varied ML outputs and subject them to rigorous, explainable logical analysis.
@@
To address these challenges, this paper introduces a novel approach that integrates the outputs of various machine learning models directly with the PyReason framework [3], an open-world temporal logic reasoning engine. Figure 1 provides a high-level overview of this integration framework. PyReason’s foundation in generalized annotated logic [8] directly incorporates real-valued outputs (e.g., probabilities, confidence scores) from diverse ML models, treating them as truth intervals within its logical framework. We note that the existing work focuses on integration between machine learning models and logic-based reasoning by either modifying the ML model (e.g., see [5, 11, 13, 6, 7, 4]) or by adding rules about the failure of the model to the logic program (e.g., see [9, 2, 10]). In this paper, we examine practical methods complementary to both approaches - focusing on software engineering questions concerning the integration between logic programs and ML models. This distinction carries a significant impact: by not requiring modifications to the ML model, our approach allows for the seamless integration of pre-trained, “black-box” systems. This enhances modularity and practical applicability, as it enables the use of state-of-the-art proprietary models or those that are computationally expensive to fine-tune, thus lowering the barrier for creating complex neuro-symbolic applications.
@@
A central pillar of our proposal is the ability of PyReason to implement logic programs with Python code. This is achieved through mechanisms that enable Python code to continuously poll ML model results, converting these perceptions into logical facts that are directly injected into the logic program. This process not only feeds the reasoning engine, but also triggers the re-computation of the minimal model, ensuring that logical decisions always reflect the latest observations from the ML model. Given the ICLP audience’s familiarity with both Python and logic programming, we have prioritized a design that explicitly intersperses both paradigms, facilitating an understanding of the interaction between ML perception and logical deduction. We have made several design decisions to simplify this integration and improve usability. For example, the introduction of specific base classes such as LogicIntegrationBase and TemporalLogicIntegratedClassifier standardizes the inference flow and the conversion of predictions into logical facts. These decisions simplify syntax, manage optional arguments, and support a robust, modular approach to complex process automation.
In addition to this integration, PyReason supports temporal reasoning, knowledge graph integration, and explainable inference traces. While traditional logic programming languages like Prolog can provide explainable reasoning paths, they are less suited for integrating existing knowledge graphs and handling time-based logic. Python, by contrast, offers better support for these needs through libraries like GraphML and NetworkX, and is more adaptable for implementing temporal reasoning. By combining outputs from machine learning models with PyReason’s logical reasoning capabilities, we aim to build a system for automating complex processes. This approach is applicable across domains such as manufacturing (using object detection for quality control), healthcare (image classification for diagnostics), and business operations (LLMs for document processing and customer support).
@@
The remainder of this paper is structured as follows. Section2 introduces the PyReason framework, its logical foundations, and its reasoning capabilities. Section3 details the integration with machine learning models via a standardized interface. Section4 presents two case studies—robotic welding and a card game—to illustrate practical applications. Section5 concludes with a discussion of the broader implications.
@@ 2 The PyReason Framework
PyReason is a software framework developed for open-world temporal logic, implemented in Python [3]. It is founded on generalized annotated logic, a system which not only encapsulates the current range of differentiable (i.e., fuzzy) first-order logics but also integrates temporal extensions to support inference over finite durations. A primary design objective of PyReason is to provide direct support for reasoning on graphical structures, such as knowledge graphs, social networks, and biological networks. The framework is engineered to generate fully explainable traces of its inference processes, ensuring transparency. It also includes type-checking and a memory-efficient design for scalable deductive inference. PyReason endeavors to supply a modular and extensible platform for deduction, addressing the constraints of bespoke software frequently developed for particular logics within neuro-symbolic frameworks. It captures a breadth of capabilities observed in diverse neuro-symbolic frameworks, encompassing fuzzy logic, open-world reasoning, temporal reasoning, and graph-based reasoning. The PyReason software is publicly accessible via GitHub (https://github.com/lab-v2/pyreason).
@@ 2.1 Core Logical Principles and Reasoning over Graphs
[…]
PyReason integrates temporal logic over finite periods through ”interpretations”, which are an initial set of facts correct at $t=0$. Temporal annotations $I(A,\hat{t}):[L,U]$ represent discrete time-points, where $\hat{t}$ is a specific time $t$ or $s$ for static interpretations (unchanging over time). Non-static predicates can change annotations over time without a monotonicity requirement. The framework operates under an open-world assumption: interpretations not explicitly defined are initially ”unknown” (e.g., [0,1][0,1][ 0 , 1 ]).
The framework is designed to reason about graphical structures, such as knowledge graphs. It assumes a graph $G=(\mathcal{C},E)$, where nodes $\mathcal{C}$ are constants, and edges $E$ specify potential relationships. Predicates are typically unary (node labels) or binary (edge labels). A reserved binary predicate ‘rel‘ exists, where $rel(a,b)$ is a tautology if $(a,b)\in E$ and uncertain otherwise. Graphical syntactic extensions allow for operators like requiring the existence of $k$ items.
@@ 2.2 Key Reasoning Capabilities and Features
Logical rules are the primary mechanism for modifying annotations of non-static predicates. PyReason supports ground rules, universally quantified non-ground rules (for reasoning within or across edges), and rules with rule-based quantifiers in the head. Rules incorporate a temporal component through a time delay, denoted as $\Delta t$. Syntactically, this delay is represented as a subscript to the implication arrow (e.g., $\leftarrow_{\Delta t}$, specifying the number of timesteps that must elapse after the body is satisfied before the head’s annotation is updated. This feature explicitly defines the temporal nature of the logic programs used within PyReason. Rule heads often involve functions $f(x_{1},...,x_{n})$ (e.g., T-norms, T-conorms, algebraic functions) to combine annotations from body literals.
@@
[…]
@@ 2.3 Implementation and Neuro-Symbolic Relevance
PyReason is a modern Python framework designed for scalability and correctness. Graphical input is via PrologML or a Networkx Prolog, with initial conditions and rules in text format. The Numba JIT compiler optimizes key operations for performance and supports CPU parallelism. For memory efficiency, only current interpretations are stored. However, past states are reconstructible via rule traces, which enable explainability for long-running implementations. This design supports canonical and non-canonical models with minimal computational overhead by recomputing interpretations at each step.
PyReason is strategically positioned as an instrument in neuro-symbolic AI. Generalized annotated logic, its foundation, captures a wide spectrum of real-valued, temporal, and fuzzy logics used in frameworks like Logical Tensor Networks (LTN) and Logical Neural Networks (LNN), both viewable as subsets of annotated logic. PyReason can conduct inference on logic programs learned or modified by such systems. It enables capabilities like graph-based and temporal reasoning, potentially absent in other neuro-symbolic frameworks’ native logics. The explainability of its inference is essential to understand and debug complex neuro-symbolic models. By supporting generalized annotated logic and its extensions, PyReason facilitates system design independent of the learning process, offering a robust deductive engine.
@@ 3 Integration with Machine Learning Models
PyReason allows users to integrate machine learning (ML) models directly within the logical framework. ML models output tensors, which are converted into PyReason facts that can be used to perform inference. This section outlines this process and illustrates its simplicity within our framework.
@@ 3.1 Logic Integrated Base Class
PyReason defines an abstract base class LogicIntegrationBase, designed for subclassing, to integrate specific machine learning models. This class declares three protected methods that are called sequentially and must be implemented by every subclass:
@@ 3.2 Logic Integrated Classifier
For simplicity, we outline the process of converting probabilities to annotated bounds as a flowchart depicted in Figure 2. By following this logic, we are able to convert model output tensors into annotated facts which can be added into PyReason for downstream inference.
@@ 3.3 Automated Temporal Logic Integrated Classifier
In many real‐world applications, data arrives continuously from sensors or cameras. Logical inferences must adapt in real time. To address this, PyReason provides the TemporalLogicIntegratedClassifier, a subclass of LogicIntegratedClassifier that ties PyReason’s internal temporal reasoning to wall‐clock time. This class adds two key parameters: poll_interval, which specifies the frequency (in seconds or PyReason time) to poll the model with new input data if the poll_condition is either satisfied or not defined; and poll_condition, a PyReason query that determines whether a new inference cycle should be triggered.
At each interval, PyReason evaluates the poll_condition. If the condition is satisfied, or if no condition is specified, the system automatically:
@@
This end‐to‐end loop (Figure 3) requires no additional glue code: PyReason handles the timing, model polling, fact injection, and inference in a single, explainable pipeline.
Section 4 illustrates two concrete workflows built on this machinery: defect detection in robotic welding and decision‐making in a simulated card game. In each case, the temporal classifier ensures that as soon as new data are available, the logic program reflects the latest observations and derives the appropriate conclusions.
@@ 4.1 Welding Classifier
In an industrial robotic welding cell, detecting weld defects in real time is essential to maintaining product quality and avoiding costly rework. Here, we apply a pre-trained weld-defect detector to identify gaps in each weld image, feed those results into PyReason to attempt automated gap closure, and mark the part as defective if the gap cannot be closed.
@@ 4.1.1 Logical Encoding in PyReason
The gap(weld_object) atom is added to the program by PyReason’s LogicIntegratedClassifier. These two rules help determine if the weld was unsuccessful and if after a repair was attempted there is still a gap, then the part is marked as defective. A flowchart of the pipeline can be seen in Figure 4. The repair and conclusion that a part might be defective run in a background routine (Thread 2) while the main thread (Thread 1) goes over each object to be welded and performs the welding task.
We use the TemporalLogicIntegratedClassifier defined in Section 3.3, with a poll_interval set to 0.5 seconds and the poll_condition set to gap(weld_object):[1,1]. This allows us to check whether there is a gap in the weld every 0.5 seconds and attempt to repair the gap in the background.
@@ 4.1.2 Interpreting the Inference Trace
Table 1 showcases an explainable PyReason trace after welding 5 objects over 7 fixed-point operations (FPO). It can be seen that both thread 1 and thread 2 update the interpretation simultaneously. Thread 1 is always called first on every new object, and thread 2 is invoked only when a repair is necessary. In FPOs 2-3, a repair is attempted due to a gap being detected in FPO 1 and the object is deemed defective after an unsuccessful repair. Similarly, another repair is invoked in FPO 5, but the repair is successful and the part is not marked as defective.
@@ 4.2 Simple Card Game
This case study demonstrates how PyReason’s open-world temporal logic can drive decision-making in a simple ”42” card game by periodically integrating the output from an image classifier and applying annotated logic to manage uncertainty over time. The objective of the game is to draw as many cards as possible without exceeding a point total of 42. Card values are defined as 3 points for an Ace, 6 for numeric cards, and 9 for face cards. At each turn, a decision is made to draw another card if there is a nonzero probability of remaining at or below the 42-point limit, based on the cards remaining in the deck. The game concludes when any subsequent draw would guarantee exceeding the limit.